# Trajectory Optimization for In-Hand Manipulation with Tactile Force Control

### v1

@@ Abstract
The strength of the human hand lies in its ability to manipulate small objects precisely and robustly. In contrast, simple robotic grippers have low dexterity and fail to handle small objects effectively. This is why many automation tasks remain unsolved by robots. This paper presents an optimization-based framework for in-hand manipulation with a robotic hand equipped with compact Magnetic Tactile Sensors (MTSs). The small form factor of the robotic hand from Shadow Robot introduces challenges in estimating the state of the object while satisfying contact constraints. To address this, we formulate a trajectory optimization problem using Nonlinear Programming (NLP) for finger movements while ensuring contact points to change along the geometry of the fingers. Using the optimized trajectory from the solver, we implement and test an open-loop controller for rolling motion. To further enhance robustness and accuracy, we introduce a force controller for the fingers and a state estimator for the object utilizing MTSs. The proposed framework is validated through comparative experiments, showing that incorporating the force control with compliance consideration improves the accuracy and robustness of the rolling motion. Rolling an object with the force controller is 30% more likely to succeed than running an open-loop controller. The demonstration video is available at https://youtu.be/6J_muL_AyE8.
@@ I Introduction
In-hand manipulation skills is in many real-world applications such as industrial assembly, which inspires the significant development of robotic motion planning strategies. Recent research in robotics has actively investigated planning of intricate, human-like hand motions [1], [2], [3]. Yet, robotic in-hand motion planning with an object still remaines challenging due to the complexity of contact dynamics, high dimensionality of the configuration space, and hardware limitations such as oversized vision-based tactile sensors [4] and fingers [5]. A robust in-hand manipulation framework was recently devised to realize a stable motion planning for a cable [6], using the 4-DOF gripper attached vision-based tactile sensor, GelSight [7]. Additionally, the framework plans finger motions based on predefined motion primitives designed manually.
Another example of elaborate motion primitives is rolling an object between two fingers, which has gained attention in robotics due to its crucial role in industry applications, such as screwing bolts [8] or inserting low-volume parts [9]. To achieve the rolling motion, the framework for learning contact dynamics with vision-based tactile feedback is proposed to roll a rigid long stick between two fingers, enhancing grasp robustness [10]. From [11], the adaptive rolling motion is achieved to the object’s width by estimating the contact area from three binarized images using two visuotactile sensors attached to a parallel gripper. In the perspective of hardware solutions, various specialized gripper designs have been proposed [12], [13], [14] to accomplish robust and efficient rolling motion. However, it is still non-trivial to achieve robust and adaptive rolling motion with high precision and success rate remains challenging, even if the motion seems straightforward for humans.
In this vein, we utilize the commercial multi-finger robotic hand that is lightweight, compact, and highly human-like. Consequently, this robotic hand can provide the most similar in-hand manipulation skills to those of humans. Additionally, a magnetic sensor-based tactile sensor that shown in Fig. 1, is lighter and more compact than visuotactile sensors [7], while still being able to detect contact forces. Finally, an efficient and straightforward trajectory optimization method can be employed to generate rolling motion primitives while ensuring a firm grasp on the object.
@@ I-A Contributions
  • Design a trajectory optimization for contact dynamics to generate multi-finger motion that continuously changes contact points along the fingers
@@ II-A In-Hand Manipulation
To perform diverse and delicate tasks required in an industrial setting, in-hand manipulation is essential. In-hand manipulation can be achieved either by dexterous hands or by utilizing external forces, which leads to extrinsic dexterity. Without complex dynamic modeling, simple primitives that consider external forces can adjust the position or orientation of objects [5]. Modeling the relationships among the external environment, the gripper, and the object allows for more accurate attainment of the desired pose of the object [15, 16, 17, 18]. Although this method takes frictional forces into account, it has limitations as it relies on accurate physical information to execute open-loop actions without predicting the actual pose of the object. Neural network methods have also been studied for planning dexterous manipulation [19, 20] but still do not estimate state of the objects that can be used for closing the control loop. [21] proposed a reinforcement learning method to learn dexterous in-hand manipulation with Shadow Dexterous hand. They successfully rotate a given object to the goal configuration. However their approach relied on multiple vision sensors, and the object was relatively large.
The limitations in pose estimation of grasped objects can be addressed with recently developed tactile sensors [7, 22, 23]. Due to the flexibility of hardware reconfiguration in parallel grippers, in-hand manipulation using tactile sensors has been extensively studied [24, 25, 26]. However, since two-finger gripper have a limited degree of the freedom, the external environment is often utilized to generate dexterous hand manipulation. Unlike parallel grippers, incorporate tactile sensors into a human-like multi-finger hand is challenging, as tactile sensors are often larger than fingertips. Therefore, various control methods have been explored to have an effective finger control for robotic hands to enhance the stable and precise in-hand manipulation.
@@ II-B Finger Control
Finger control can be achieved in several ways, tailored to the specific control objectives. The impedance control scheme in torque-controlled robotic fingers is employed in [27] and extended with learning-based control parameter estimation to enhance the performance of the controller [28]. The finger dynamics model can be used to implement inverse dynamics control for finger control [29], based on a joint torque controller. Besides these control methods, hybrid position/force control [30], force-position control [31], force-velocity control [32] and position-velocity control [33] provide alternative control designs for fingers to deal with intricate interactions between fingers and the contact environments. The choice of control strategy depends on specific task requirements and different robot hand systems.
@@ III-A Contact Dynamics for Rolling Motion
Consider two fingers in contact with an object as shown in Fig. 3. The forces exerted onto the object at the contact points $A$ and $B$ are called $\mathbf{F}_{AO}$ and $\mathbf{F}_{BO}$. The following describes the quasi-static dynamics of the system similar to [34]. To enable rolling, sustained contact must be obtained between both fingers and the object. Therefore, we ensure that the contact dynamics holds the quasi-static equilibrium of the object between the fingers as follows:
@@
where $\|\bm{f}_{t}^{i}\|$ denotes the 2-norm of $\bm{f}_{t}^{i}$ and at contact point $i\in\{A,B\}$, $\mu_{i}$ is the coefficient of friction between the fingertip and the object. By approximating the contact as a point contact, moment friction is neglected.
Rolling an object requires a continuous transition of contact points while maintaining persistent contact. Contact points A and B must transition along the fingers to achieve proper rolling. However, contact-rich manipulation tasks are usually performed with stationary contact. To generate a trajectory that allows contact point B to change, we add a virtual prismatic joint to the finger. As shown in Fig. 3, $d(t)$ is a time-dependent variable that corresponds to the length of the virtual prismatic link, which changes as the object rolls between the fingers to maintain contact. Contact between the virtual prismatic link and the object is enforced by:
$d_{k-1}-d_{k}-r(\theta_{o,k-1}-\theta_{o,k})=0$ (4)
where $\theta_{O}$ represents the orientation of the object and $k$ denotes a time step for the trajectory and $r$ is radius of object. Under the no-slip assumption, the state of the object can be represented in a manner similar to (4).
$\|\mathbf{c}_{k-1}-\mathbf{c}_{k}\|-r(\theta_{o,k-1}-\theta_{o})=0$ (5)
@@ III-B Offline Motion Planning
Trajectory optimization is used to plan an offline rolling motion reference for the first finger. To simplify the problem, the thumb is considered flat as the possible travel distance for the object is relatively small. The contact model is formulated as a complementarity problem alongside trajectory optimization, incorporating complementarity constraints as presented in [35]. We assume that the object remains in quasi-static equilibrium, considering the simplified planar dynamics. The optimization is performed in discrete time, with a horizon of $N$ time steps. The trajectory optimization for rolling an object is formulated as follows:
@@
The term $\psi$, which represents the distance between the fingers and the object, is added to the cost function to enforce contact between the fingers and the object. Incorporate $\psi$ in the cost function is crucial as it ensures that fingers reach the object and apply force while maintaining contact. As shown in Fig. 3, the initial configuration of the first finger is not in contact with the object.
Another important constraint is (4), which defines a virtual prismatic joint that enables rolling while allowing the contact point between the finger and the object to change continuously. If constraint (4) is not included in the optimization problem, the solver is likely to fail to find a solution. In such case, the finger may make contact only at the fingertip, resulting in the object being pushed by the fingertip rather than naturally following the geometry of the fingers during rolling. This nonlinear problem (6) was solved using IPOPT [36] in CasADi [37]. The solution consists of fingertip positions and force information, which serve as the reference trajectory for the rolling motion.
@@ III-C1 Contact Point Estimation
Let $\bm{f}_{i},\bm{p}_{i}\in\mathbb{R}^{3}$ be the magnetic sensor values and the sensor positions at each finger, respectively, where $i\in I=\{1,\dots,n_{s}\}$ and $n_{s}$ represents the total number of sensors attached to the fingers. Define $S\in\mathbb{R}^{n_{s}}$ as a vector with elements $S_{i}=||\bm{f}_{i}||$, where each $S_{i}$ corresponds to $\bm{p}_{i}$ and its magnitude represents $||\bm{f}_{i}||$. The vector $S$ is normalized to a range from 00 to $255$ as follows:
$S^{{}^{\prime}}_{i}=255\bigg{(}\frac{S_{i}-\min(S)}{\max(S)-\min(S)}\bigg{)},$ (7)
@@
$W_{i}=\max\left(\frac{S^{{}^{\prime}}_{i}-T}{S^{*}-T},0\right),$ (8)
Let $J\subseteq I\setminus\{i_{*}\}$ be a set of indices corresponding to the $n_{n}$, nearest neighbor, elements such that
$\|p_{i^{*}}-p_{i}\|\leq\|p_{i^{*}}-p_{j}\|\quad\forall(i,j)\in J\times I\setminus J^{-}$ (9)
Finally, the contact point is computed as:
$\bm{p}_{\textnormal{cp}}=\sum_{i\in J}\tilde{w}_{i}\bm{p}_{i}\textnormal{ ,}\quad\tilde{w}_{i}=\frac{S^{{}^{\prime}}_{i}}{\sum_{j\in J}{S^{{}^{\prime}}_{j}}},$ (10)
@@ III-D Finger Motion and Force Control
However, tracking a reference trajectory using only position and velocity control is insufficient for stable object manipulation. Since the manipulation task is simplified for trajectory optimization, an open-loop controller may fail to maintain a stable grasp during the rolling motion. While the intrinsic compliance of the hand aids stability, it cannot fully compensate for significant force variation. To address this issue, we incorporate a force controller to ensure robust manipulation. Additionally, a compliance analysis was conducted to account for finger deformations during the force control process.
@@ III-D1 Finger Controller
where $\bm{p}_{\textnormal{ref}}\in\mathbb{R}^{3}$ and $\bm{p}_{\textnormal{c}}\in\mathbb{R}^{3}$ are the reference and current measured positions of a fingertip, respectively. The position of the tip can be obtained from forward kinematics of the hand. Subsequently, we use proportional–integral–derivative (PID) controller to define the reference velocity $\bm{v}_{\textnormal{ref}}\in\mathbb{R}^{3}$
In the same way, the PID controller scheme is also utilized to track $\bm{v}_{\textnormal{ref}}$ with the current $\bm{v}_{c}\in\mathbb{R}^{3}$ defined as:
@@
$\Delta\dot{\bm{q}}=J^{+}(\bm{q})\Big{(}\underbrace{k_{\textrm{pv}}\tilde{\bm{v}}_{\textnormal{e}}+k_{\textrm{iv}}\int\tilde{\bm{v}}_{\textnormal{e}}\textnormal{dt}+k_{\textrm{dv}}\dot{\tilde{\bm{v}}}_{\textnormal{e}}}_{\tilde{\bm{v}}}\Big{)},$ (14)
where $\tilde{\bm{v}}_{\textnormal{e}}=\bm{v}_{\textnormal{ref}}-\bm{v}_{\textnormal{c}}$, and $k_{\textrm{pv}},k_{\textrm{iv}},k_{\textrm{dv}}\in\mathbb{R}$ are PID gains for the velocity controller. However, in (14), $\Delta\dot{\bm{q}}$ is determined based on the pseudo-inverse matrix, which may cause $\bm{q}_{\textnormal{ref}}$ to deviate significantly from $\bm{q}_{\textnormal{c}}$, leading to discontinuous controller behavior. To resolve this, we enforce a constraint ensuring that $\bm{q}_{\textnormal{ref}}$ remain close to the current configuration of each finger. Consequently, we formulate the following linear least-squares problem:
$\min_{\Delta\dot{\bm{q}}}\frac{1}{2}{||\tilde{\bm{v}}-J(\bm{q})\Delta\dot{\bm{q}}||}^{2}$ (15)
@@ III-D2 Force Controller
Kinematic model and simulations of the Shadow Hand do exist [38]; however, this model does not include the coupling between individual finger joints and their compliance when external forces are applied. Consequently, we provide such a model and show how to integrate the coupling mode into the force controller.
Figure 6: (a) Changes in $q_{1}$ when $\tau$ is applied. (b) Compliance of Joint 1 of the first finger of the Shadow Robot Hand.
Fig. 6(a) shows the $q_{1}$ of the first finger when a force is applied at the fingertip, while Fig. 6(b) indicates the variation in compliance depending on the reference angle for $q_{1}+q_{2}$. It is seen that the compliance increases linearly from 00 to $\frac{\pi}{2}$ and then decreases linearly after $\frac{\pi}{2}$. This trend implies that the mechanical structure of the finger exhibits variable compliance, which should be considered when performing force control with the finger.

### v2

@@ Abstract
The strength of the human hand lies in its ability to manipulate objects precisely and robustly. In contrast, simple robotic grippers have low dexterity and fail to handle small objects effectively. This is why many automation tasks remain unsolved by robots. This paper presents an optimization-based framework for in-hand manipulation with a robotic hand equipped with compact Magnetic Tactile Sensors (MTSs). We formulate a trajectory optimization problem using Nonlinear Programming (NLP) for finger movements while ensuring contact points to change along the geometry of the fingers. Using the optimized trajectory from the solver, we implement and test an open-loop controller for rolling motion. To further enhance robustness and accuracy, we introduce a force controller for the fingers and a state estimator for the object utilizing MTSs. The proposed framework is validated through comparative experiments, showing that incorporating the force control with compliance consideration improves the accuracy and robustness of the rolling motion. Rolling an object with the force controller is 30% more likely to succeed than running an open-loop controller. The demonstration video is available at https://youtu.be/6J_muL_AyE8.
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
In-hand manipulation skills are in many real-world applications such as industrial assembly, which inspires the significant development of robotic motion planning strategies. Recent research in robotics has actively investigated planning of intricate, human-like hand motions [Jihong2020], [wilson2023cable], [Manes2024]. Yet, robotic in-hand motion planning with an object still remains challenging due to the complexity of contact dynamics, high dimensionality of the configuration space, and hardware limitations such as bulky vision-based tactile sensors [she2021cable] and fingers [Nikhil2014SimpleGripper].
Among various motion primitives, rolling an object between two fingers serves as a representative example of complex in-hand manipulation. This primitive has gained attention in robotics due to its crucial role in industry applications, such as screwing bolts [Tang2024Screw] or inserting low-volume parts [Azulay2023RAL]. To achieve the rolling motion, the framework for learning contact dynamics with vision-based tactile feedback is proposed to roll a rigid long stick between two fingers, enhancing grasp robustness [du2024stick]. From [Zhang2022Rolling], the adaptive rolling motion is achieved to the object’s width by estimating the contact area from three binarized images using two visuotactile sensors attached to a parallel gripper. In the perspective of hardware solutions, various specialized gripper designs have been proposed [Chen2014IROS], [Zuo2021], [Yan2022ICMA] to accomplish robust and efficient rolling motion. However, it is still non-trivial to achieve robust and adaptive rolling motion with high precision and success rate, even if the motion seems straightforward for humans.
In this vein, we utilize a commercial multi-finger robotic hand that is lightweight, compact, and highly human-like. Consequently, this robotic hand can provide the most similar in-hand manipulation skills to those of humans. Additionally, a magnetic tactile sensor shown in Fig. 1, is lighter and more compact than visuotactile sensors [yuan2017gelsight], while still being able to detect contact forces. Finally, an efficient and straightforward trajectory optimization method can be employed to generate rolling motion primitives while ensuring a firm grasp on the object.
@@ I-A Contributions
  • Design a trajectory optimization for contact dynamics to generate multi-finger motion that continuously changes contact points along the fingers.
@@ II-A In-Hand Manipulation
To perform diverse and delicate tasks, in-hand manipulation is essential. In-hand manipulation can be achieved either by dexterous hands or by utilizing external forces, which leads to extrinsic dexterity. Without complex dynamic modeling, simple primitives that consider external forces can adjust the position or orientation of objects [Nikhil2014SimpleGripper]. Modeling the relationships among the external environment, the gripper, and the object allows for more accurate attainment of the desired pose of the object [mordatch2012contact, chavan2015prehensile, chavan2020planar, hou2018fast]. While frictional forces are taken into account, the method remains limited as it execute open-loop actions without predicting the actual pose of the object. Learning-based methods have also been studied for planning dexterous manipulation [nagabandi2020deep, kumar2016optimal] but still do not estimate state of the objects that can be used for closing the control loop. [andrychowicz2020learning] proposed a reinforcement learning method to learn dexterous in-hand manipulation with Shadow Dexterous hand. They successfully rotate a given object to the goal configuration. However their approach relied on multiple vision sensors, and the object was relatively large.
The limitations in pose estimation of grasped objects can be addressed with recently developed tactile sensors [yuan2017gelsight, lambeta2020digit, taylor2022gelslim]. Due to the flexibility of hardware reconfiguration in parallel grippers, in-hand manipulation using tactile sensors has been extensively studied [hogan2020tactile, kim2023simultaneous, bronars2024texterity]. However, since two-finger gripper have a limited degree of the freedom, the external environment is often utilized to generate dexterous hand manipulation. Incorporating tactile sensors into a human-like multi-finger hand is challenging, as tactile sensors are often larger than fingertips. Therefore, various control methods have been explored to have an effective finger control for robotic hands to enhance the stable and precise in-hand manipulation.
@@ II-B Finger Control
Finger control can be achieved in several ways, tailored to the specific control objectives. The impedance control scheme in torque-controlled robotic fingers is employed in [ButterfassDLRHand2001] and extended with learning-based control parameter estimation to enhance the performance of the controller [Sievers2022]. The finger dynamics model can be used to implement inverse dynamics control for finger control [Xue2023IROS], based on a joint torque controller. Besides these control methods, hybrid position/force control [YOSHIKAWA2010199], force-position control [ROJASGARCIA2022233], force-velocity control [Li-RSS-13] and position-velocity control [Raković2018] provide alternative control designs for fingers to deal with intricate interactions between fingers and the contact environments.
@@ III-A Contact Dynamics for Rolling Motion
Consider two fingers in contact with an object as shown in Fig. 3. The forces exerted onto the object at the contact points $A$ and $B$ are called $\mathbf{F}_{AO}$ and $\mathbf{F}_{BO}$. The following describes the quasi-static dynamics of the system similar to [shirai2023tactile]. To enable rolling, sustained contact must be obtained between both fingers and the object. Therefore, we ensure that the contact dynamics holds the quasi-static equilibrium of the object between the fingers as follows:
@@
where $\|\bm{f}_{t}^{i}\|$ denotes the 2-norm of $\bm{f}_{t}^{i}$ at contact point $i\in\{A,B\}$ and $\mu_{i}$ is the coefficient of friction between the fingertip and the object. By approximating the contact as a point contact, moment friction is neglected.
Rolling an object requires a continuous transition of contact points while maintaining persistent contact. Contact points A and B must transition along the fingers to achieve proper rolling. However, contact-rich manipulation tasks are usually performed with stationary contact. To generate a trajectory that allows contact point B to change, we add a virtual prismatic joint to Finger B. As shown in Fig. 3, $d(t)$ corresponds to the length of the virtual prismatic link, which changes as the object rolls between the fingers to maintain contact. Contact between the virtual prismatic link and the object is enforced by:
$d_{k-1}-d_{k}-r\cdot(\theta_{o,k-1}-\theta_{o,k})=0$ (4)
where $\theta_{O}$ represents the orientation of the object and $k$ denotes a time step for the trajectory and $r$ is the radius of object. Under the no-slip assumption, the state of the object can be represented in a manner similar to (4).
$\|\mathbf{c}_{k-1}-\mathbf{c}_{k}\|-r\cdot(\theta_{o,k-1}-\theta_{o})=0$ (5)
@@ III-B Offline Motion Planning
Trajectory optimization is used to plan an offline rolling motion reference for the first finger. To simplify the problem, the thumb is considered flat as the possible travel distance for the object is relatively small. The contact model is formulated as a complementarity problem alongside trajectory optimization, incorporating complementarity constraints as presented in [posa2014direct]. We assume that the object remains in quasi-static equilibrium, considering the simplified planar dynamics. The optimization is performed in discrete time, with a horizon of $N$ time steps. The trajectory optimization for rolling an object is formulated as follows:
@@
The term $\psi_{k}$ represents the distance between the fingers and the object, is added to the cost function to enforce contact between the fingers and the object. Incorporation of $\psi$ in the cost function is crucial as it ensures that fingers reach the object and apply force while maintaining contact. As shown in Fig. 3, the initial configuration of the first finger is not in contact with the object.
Another important constraint is (4), which defines a virtual prismatic joint that enables rolling while allowing the contact point between the finger and the object to change continuously. If constraint (4) is not included in the optimization problem, the solver is likely to fail to find a solution. In such case, the finger may make contact only at the fingertip, resulting in the object being pushed by the fingertip rather than naturally following the geometry of the fingers during rolling. The nonlinear problem (6) was solved using IPOPT [wachter2006implementation] in CasADi [andersson2019casadi]. The solution consists of fingertip positions and force information, which serve as the reference trajectory for the rolling motion.
@@ III-C1 Contact Point Estimation
[…]
$s^{{}^{\prime}}_{i}=255\bigg(\frac{s_{i}-\min(\textbf{s})}{\max(\textbf{s})-\min(\textbf{s})}\bigg),$ (7)
@@
$W_{i}=\max\left(\frac{s^{{}^{\prime}}_{i}-T}{s_{i^{*}}^{{}^{\prime}}-T},0\right),$ (8)
Let $J\subseteq I$ be a set of indices corresponding to the $n_{n}$ nearest neighbors elements. Then, the contact point is computed as:
$\bm{p}_{\textnormal{cp}}=\sum_{i\in J}\tilde{w}_{i}\bm{p}_{i}\textnormal{ ,}\quad\tilde{w}_{i}=\frac{s^{{}^{\prime}}_{i}}{\sum_{j\in J}{s^{{}^{\prime}}_{j}}}$ (9)
@@ III-D Finger Motion and Force Control
However, tracking a reference trajectory using only position and velocity control is insufficient for stable object manipulation. Since the manipulation task is simplified for trajectory optimization, an open-loop controller may fail to maintain a stable grasp during the rolling motion. While the intrinsic compliance of the hand aids stability, it cannot fully compensate for significant force variation. To address this issue, we incorporate a force controller to ensure robust manipulation. Additionally, a compliance analysis of tendons was conducted to account for finger deformations during the force control process.
@@ III-D1 Finger Controller
where $\bm{p}_{\textnormal{ref}}\in\mathbb{R}^{3}$ and $\bm{p}_{\textnormal{c}}\in\mathbb{R}^{3}$ are the reference and current measured positions of a fingertip, respectively. The position of the tip can be obtained from forward kinematics of the hand. Subsequently, we use a proportional–integral–derivative (PID) controller to define the reference velocity $\bm{v}_{\textnormal{ref}}\in\mathbb{R}^{3}$
In the same way, a PID controller is utilized to track $\bm{v}_{\textnormal{ref}}$ with the current $\bm{v}_{c}\in\mathbb{R}^{3}$ defined as:
@@
[…]
where $\tilde{\bm{v}}_{\textnormal{e}}=\bm{v}_{\textnormal{ref}}-\bm{v}_{\textnormal{c}}$, and $k_{\textrm{pv}},k_{\textrm{iv}},k_{\textrm{dv}}\in\mathbb{R}$ are PID gains for the velocity controller. In (13), $\Delta\dot{\bm{q}}$ is determined based on the pseudo-inverse matrix, which may cause $\bm{q}_{\textnormal{ref}}$ to deviate significantly from $\bm{q}_{\textnormal{c}}$, leading to large changes in the controller output. To resolve this, we enforce a constraint ensuring that $\bm{q}_{\textnormal{ref}}$ remain close to the current configuration of each finger. Consequently, we formulate the following linear least-squares problem:
[…]
@@ III-D2 Force Controller
Kinematic model and simulations of the Shadow Hand do exist [menagerie2022github]; however, this model does not include the coupling between individual finger joints and their compliance when external forces are applied. Consequently, we provide such a model and show how to integrate the coupling mode into the force controller.
Figure 6: (a) Change in $q_{1}$ during a single trial with varying external torque $\tau$. (b) Compliance of the first finger under various finger configurations.
Fig. 6(a) shows $q_{1}$ of the first finger when a varying force is applied at the fingertip, while Fig. 6(b) indicates the variation in compliance depending on the reference angle for $q_{1}+q_{2}$. It is seen that the compliance increases linearly from 0 to $\frac{\pi}{2}$ and then decreases linearly after $\frac{\pi}{2}$. This trend implies that the mechanical structure of the finger exhibits variable compliance, which should be considered when performing force control with the finger.